# Effective Multi-Dimensional 3D Scatterplots as 2D Figures

### v2

@@ Abstract
Computationally and data intensive workloads including design space exploration or large studies often lead to multi-dimensional results, which are often not trivial to digest with conventional plotting software. 3D scatterplots can be a powerful technique to visualise and explore such datasets, especially with the help of colour mapping and other approaches to represent more than the 3 dimensions of the Cartesian coordinate system. However, modern software commonly lacks this multi-dimensional functionality or is ineffective. One limitation is the frequent use of isometric axes, which is equivalent to removing one entire dimension. In manuscripts, additional visual cues such as movement are also not present to mitigate for the loss of depth perception and spatial information, hence their relatively limited use as static figures.
In this work, we present a novel open-source JavaFX-based plotting framework that focuses on easy exploration of multi-dimensional datasets, and provides unique features or feature combinations to improve knowledge transfer from single stand-alone plots. An empirical study was conducted within an academic institution to quantify the effectiveness of feature or feature combinations on 3D scatterplots in terms of reading time and accuracy.
Index Terms:
3D scatterplots, multi-dimensional, design space, visual cues, depth perception, discrete data, empirical study
I Introduction
@@
Effectively visualising and analysing multi-dimensional data is of prime importance for a variety of sciences, and is an invaluable tool in data mining, machine learning and statistics. With global trends and events such as large-language models [1] and pandemics [2], intuitively visualising large or complex data are already concerning a wider demographic. A visualisation method that is appropriate for multi-dimensional data is 3D scatterplots, as they can translate data for three or more dimensions into the 3D space, but additional care is needed to produce effective and informative plots.
@@
When compared to 2D scatterplots, having an extra dimension in space can be a blessing and a curse. It is a blessing because a single 3D scatterplot can give more advanced insights including how each of the dimensions correlate with each other or collectively. When combined with additional cues such as the point size, 3D scatterplots can also provide quicker insights for complex problems, as well as when the data can be explained effectively with spatial information. It is also a curse, however, because in their static form as in scientific manuscripts, they need to be flattened to fit in two-dimensional means. These essentially reduce to 2D projections of (pseudo) 3D models. Such projections usually lack realism and visual cues to help with depth perception, for example, and this introduces ambiguity in the readings.
Due to the limitations of common graphing software and research, 3D scatterplots are only usually found in specialised applications [3]. In scientific paper publications and technical reports, figures are by nature static and two-dimensional, and can magnify the shortcomings of existing 3D scatterplotting techniques. The absence of any reader interaction from static 3D scatterplots makes it more challenging to communicate any depth information, and often limits the options for data visualisation to two-dimensional plots. The functionality of 3D scatterplots or multiple dimensions is also not a given in software with a low barrier to entry including office suites.
@@
This paper provides a novel open-source plotting framework and explores the hypothesis that the use of certain visual cues can significantly help with the readability and effectiveness of 3D scatterplots. The explored visual cues include the field of view (FoV), fog, depth-based colour mapping, projections, light paths and surface ticks. We illustrate our approach using the framework, and evaluate the effectiveness of these visual cues with a study with 57 participants from an academic community. The study measures the time and accuracy of the user measurements for random data to identify impactful visual cue combinations and settings.
The limitations in existing software and relevant research on 3D scatterplots relate to the lack of realism of the produced plots, and the ability to represent and communicate multiple-dimensions. A frequent constraint is the use of isometric projection, with distant objects having the same size as close objects, or close to isometric axes with a field of view near 0. This is equivalent to removing an entire dimension. A related issue is the use of raster or static figures as points and labels, which also lack the expected FoV-related transformations as would be seen in the 3D space.
@@
Motivation: Figure 1 shows 3D scatterplots from popular open source plotting software (Gnuplot and GNU Octave) that demonstrate common weaknesses. The default configurations (without additional scripting or code) yield same-sized points, isometric axes, and absence of additional indicators to track the coordinates of the points. This experiment uses a random distribution dataset as a worse case (as in section IV-A), but the discrete nature of the data is easily relatable to existing problems, such as in computer architecture [4]. These plots can be considered counterexamples for effectively plotting 3D scatterplots, though more complex design decisions and software are capable of adding more realism, but with different shortcomings (see section V-A).
@@
Figure 1: Default 3D plots from modern software on an arbitrary dataset.
Contributions: The main contribution is the presented open-source plotting framework [^1] Available after publication.. It provides novel visual cues or visual cue combinations for effectively communicating depth information and multiple dimensions. In contrast to related work, this software highly focuses on quickly and effectively expressing multiple-dimensions with realistic 3D scatterplots, such as with the use of unique features.
The list of unique features can be organised according to their corresponding aims, and is as follows:
  • For improving depth perception: depth-based colour mapping (traditionally only for an axis), fog emulation, a highly adjustable FoV.
  • For increasing the readability of the plotted points and labels in the 3D space: multi-axis light paths (alternative to direction drop lines), light paths as 3D objects shaped according to the points (instead of same-sized overlaid lines), 2D projections corresponding to the point shape and colour (not disassociated), 2D surfaces as grids, labels as 3D objects (no overlays, while still facing the camera).
  • For aiding the exploration of multiple dimensions: association panel with a high number of possible dimensions, automatic discrete data support such as with strings, automated label selection though sorting an additional dimension.
Section II describes the above in detail, and section V-A discusses their standing amongst state-of-the-art frameworks.
@@
Additionally, a systematic empirical user study is conducted that evaluates the visual cues in a quantifiable way in terms of measurement error (accuracy) and time (also relating to intuitiveness). The visual cues are organised in question types (see section IV-A). The outcomes of this study can be used to increase the effectiveness of 3D scatterplots and plotting software by introducing any missing beneficial features, or enabling them by default. This addresses the absence of such a quantitative study in research, which can be the reason for software of highly-varying quality in 3D scatterplots, and their relatively limited use on paper.
Following is the proposed framework (section II), example use cases (section III), the empirical study (section IV), a literature review on the related frameworks and research (section V), and finally a discussion (section VI) on insights and limitations, and a conclusion (section VII).
II Proposed framework
The presented visualisation framework generates easily-readable 3D scatterplots for multi-dimensional data, and is written in JavaFX. This functionality of the framework is controlled by a graphical-user interface (GUI). The features (sections II-A and II-B) are introduced with the help of a screenshot of the control window (figures 2 and 4 (left)), after which section II-C elaborates on the system architecture.
II-A Multiple-dimensions
The main feature of the software is the ability to effectively represent multi-dimensional data. Apart from the 3 dimensions from the Cartesian coordinate system ($x$, $y$ and $z$), the software also supports colour mapping, point size, point transparency and label overlay (namely $z1$ to $z4$ respectively). These additional dimensions are naturally supported in a one-to-many relationship between the available variables (column names) from tabular data. This arrangement currently totals 7 dimensions.
Illustratively, figure 4 (described in section III-A) shows how columns from a dataset can be assigned to any of the available dimensions in the same way as it is done for $x$, $y$ and $z$. The column labels from the dataset are transposed (becoming rows) and used here as rows of checkboxes to facilitate the assignment selection by the user. The column “Variable name” contains the names of the original columns in the dataset. $x$, $y$ and $z$ represent the corresponding axes, but are also colour-coded to ease intuitiveness. In existing plotting software the logarithmic scale is usually associated with an axis, but here there is a checkbox for each dimension to enable a quick and interactive exploration of multi-dimensional datasets.
Whenever the data of a dimension is not numerical, a dictionary (hashmap) is automatically built to represent unique strings with an integer, starting with 0 for the first occurrence. Whenever applicable, the corresponding string appears in each tick mark as opposed to its unique identifier for the first occurrence in the dictionary. This extends the applicability of the framework, such as to still be able to place person IDs on the axes or points from anonymised medical data. This functionality is absent from common plotting frameworks such as Gnuplot and GNU Octave, as the programmer has to hard-code the tick associations with strings for every axis when desired.
Labels can be placed on each point, and based on the dimension assignment they can serve as an indicator from an existing dimension, or represent a different variable. This is particularly useful when a unique string is associated with each point, as with names for instance, where placing them across one of the axes would be less helpful (having little reuse or spatial locality). For practicality reasons, the software can involve another dimension to make a subset selection. Since having a label per point can result in crowding or occlusion, the user can select the top or bottom few labels based on the values of another dimension. This exists as a layered pane inside the control window, and is unfrozen whenever labels ($z4$) are enabled (as shown on the bottom left of figure 4). See section III-B and its figure 5 for an example use case.
II-B Depth perception and point readability
One of the highlights of the presented software is the focus on depth perception. At the bottom of the window in figure 2, a layered pane includes the depth-perception-related visual cues, as elaborated below:
Figure 2: Adding visual cues through a comprehensive “View” options tab.
  • a)
    Field-of-View (FoV): Exaggerating the FoV can improve the readability of 3D scatterplots, such as by influencing the point geometry as seen by the camera (observer). The value starts from a value of 3 to avoid some JavaFX-related artifacts from overlapping surfaces.
  • b)
    Depth-based colour mapping: This feature replaces the $z1$ dimension (colour map) with the values from calculating the distance of each point to the camera. The colour palette is viridis that is designed to help with readability [5]. The corresponding variation in the colour of each point is also applied on their projection, and this is expected to help with associating arbitrary points to their projections.
  • c)
    Fog: Similar to the above, fog is designed to help with depth perception as it impacts the visibility of the points. See figure 7D for an example plot with the fog effect.
  • d)
    Projections and light paths: Projections are just two-dimensional projections on the surfaces of the view box. They inherit the colour mapping and shape of the points (e.g. circles for spheres), and there is the option to enable them in the $xz$, $yz$ and $xy$ panes (2-select combinations of the axes). The light paths connect the projection perimeters to the actual points, similar to the umbra in shadows in the real world. Projection and light path examples can be seen in figure 4.
  • e)
    Grid options: The framework supports two types of grids. First is the line, which is more conventional. Second is the grid surfaces, which can be seen as semi-transparent 3D equivalents of the grid lines, also corresponding to the tics. The grid surface is demonstrated in the example of figure 7I.
The “type” combobox for the latter two dictates the functionality of the $x$, $y$ and $z$, or $xz$, $yz$ and $xy$ checkboxes. In the shadow case, “projection” only enables the corresponding protections, as with “light path” for light paths, and “both” synchronises the appearance of both per axis. “Exclusive” falls back to “projection”, but enables a light path where a projection is available, and “preselect” operates as “light path” while enabling all projections. The idea behind the” exclusive” option is that the information of both types of shadows is redundant, but the other options exist such as when a crowded plot is in consideration. Similarly for the grid case, it is used to select between lines or surfaces, with the general idea to compact most-likely useful functionality in the GUI.
There are also features revolving around the readability of points and the legibility of the labels and tick marks. The heading and roll of the labels and their optional borders is automatically adjusted to always face the observer, while still interacting with the other cues of the 3D space including the size. The remaining customisation options are summarised in table I.
Table I: Summary of secondary customisation options
@@ II-B Depth perception and point readability
¿c c ¿c c ¿l c ¿l[colortbl-like] Component&RangesDescriptionOther notes
Frictionless spin {true, false} steady yaw rotation after cursor modification infinite animation effect, relating to depth perception
Point size 0 to 100% sizing of the points and the corresponding shadows also effective when point size is data-dependant
Point type {box, sphere, bar} selecting a point shapecan affect FoV perception
[…]
Opacity 0 to 100% $\times$ 2 adjust opacity, (influences visibility in denser data) one is for all grid types, the other is for all shadow types
Label text(string)×$\times 3$ setting the $x,y$ and $z$ axis labels pre-loaded with column names, when available
Label/tick distance -100 to 100% ×$\times 3\times 2$ relative distance from the corresponding axis bar 6 spinners for 3 axes times 2 for the labels and tick marks
Label/tick show {true, false} ×$\times 3\times 2$ enabling the labels as above
Horizontal {true, false} ×$\times 3\times 2$ rotating the corresponding labels by multiples of 90 yaw can automatically adjust increments to face the center
Synchronise {true, false} ×$\times 2$ freeze $y$ and $z$ label/tick options and mirror $x$ values simplifies controls
Borders {true, false} drawing a grey border around each label/ tick aiming at legibility
Label size 1 to 100 ×$\times 2$ label font size one for label and ticks, one for the labels of $z4$ dimension
II-C System architecture
The core functionality of the code is divided into two main Java files, one for the controls form, and the one for the interactive 3D plot. The first is mostly a Java Swing (GUI component library that comes with JDK) piece of software utilising a comprehensive selection of 2D graphical elements for interfacing with the user as a desktop application. It has references to JavaFX-related classes to be able to alter the functionality of the other class. The second file describes the JavaFX stage and is the window where the plots appear.
Figure 3 presents a high-level overview of the class instantiations inside the stage class. This representation is reminiscent of UML (unified modelling language) diagrams, but only focuses on a selection of objects that have a spatial connotation here. As can also be observed by the figure, the stage class is responsible for initiating a series of XForms in a hierarchy to effectively manipulate the plots, while avoiding code duplication. For instance, the tics for an axis are children of the same parent Xform in order to be able to hide and show them with a single high-level function call (setVisble(), inherited from the Node abstract class).
Figure 3: High-level overview of the class instantiations inside the primary stage of our JavaFX codebase.
Figure 4: Assigning columns to dimensions from a .csv dataset for a 5D plot. The dataset describes the top 500 supercomputers [6].
A high-level summary of the system architecture goes as follows. The primary stage is the window form containing the JavaFX scene, which in turn features the root node. The root node initiates the perspective camera through a series of nested Xforms for easily manipulating the point of view in the 3D space (such as for rotating, panning and zooming). The world is a child of the root node and contains all physical objects. The left cluster of objects in the diagram (figure 3) is everything related to the plotted points. There are Xforms for the point array (either boxes or spheres), projections (squares or circles, according to the point type), and similarly, light paths (boxes or cylinders). The center cluster of objects relates to the axes organisation, including the axes themselves (cylinders), secondary axes (“mirrors” in the view tab), and FXyz3D text meshes and label borders for both the axes and ticks. The last cluster of world children relate to grid lines and surfaces, and similar to ticks, they are organised internally into an array or group of Xforms, one for each axis. Their cardinality is sometimes more than 3 to facilitate the rotation functionality. For example, if an array of grids for the $z$ axis is spun to the back of the figure, they are automatically hidden, and their mirrored equivalents appear in another edge of the view box. Finally, as with the grid surfaces, the fog is emulated using a series of parallel semi-transparent surfaces, but they always face the camera and are only placed within the range of the data points.
Software dependencies: The framework is written in JavaFX, which is a Java-based library that includes three dimensional graphing capabilities, such as for building video games and visualisations. JavaFX was originally released by Sun Microsystems in 2008, and was being included in the Java Development Kit (JDK) up until Oracle corporation’s version 10 in 2018. JavaFX is still actively maintained by Oracle, but released on a separate software package licensed under the GPL-2.0 open-source license. There is no stand-alone reason behind the selection of JavaFX other than being a good general fit to our practices. It did provide a stable, mature and open platform for such a research-oriented framework working as a traditional and Unix-friendly desktop application.
Another dependency is FXyz3D, a “JavaFX 3D Visualization and Component Library” [7]. This allows dynamic creation and manipulation of text meshes in the 3D space. The idea is for the axis labels and tick labels to always face the camera, giving the impression of easily-legible overlaid text, while benefiting by having three dimensional attributes (e.g. size responding to the FoV for tick readability). Otherwise, JavaFX Text labels would only produce raster text, which even with the caching facility proved blurry for our use. The resulting software is last tested on a Windows and Arch machine with Java JDK 22 and OpenJFX 21, and uses Apache Maven for deployment.
III Use cases
This section presents example use cases demonstrating how a single plot can promptly give insights from multi-dimensional data.
III-A Top 500 supercomputers
This dataset comes from Top500.org that provides a list of the top 500 supercomputers in the world, alongside a comprehensive list of characteristics for each supercomputer [6]. The dataset can be downloaded from the website in a spreadsheet format, which can easily be parsed into a .csv (comma separated value) file, and opened directly by the framework. The generated plot is shown in figure 4 (right).
Based on the dimension assignment, as shown in figure 4 (left), the $x$-axis (red) is the rank of the supercomputer entry. The $y$-axis (blue) is the total number of CPU (central processing unit) cores in the supercomputer, and the $z$-axis is the total power in kilowatt-hours. The latter-two have their $log$ checkbox checked for applying a logarithmic scale, due to the high variability in the orders of magnitudes for those values. The $z1$ dimension is the colour and is applied to the Accelerator/Co-processor variable. This essentially means that a purple sphere denotes a CPU-only supercomputer, whereas a yellow sphere represents a supercomputer with accelerators (mainly graphics-processing units (GPUs)). Finally, the $z2$ dimension is assigned to the “Cores per Socket” variable that is an interesting attribute for the processors, and relates to the internal organisation of the “Total Cores” in processors (CPU packages).
There are numerous insights that can be derived by looking at the five-dimensional plot of figure 4 (right). First, if we ignore the 4th and 5th dimensions, there seems to be a high correlation between the rank and the total number of cores and power, when we consider them in pairs. If we take the projections on the $xy$ surface (the floor), the more performant (“faster”) the supercomputer is (rank), the more likely it is to have a higher number of cores. Similarly, the higher the core count, the more likely it is to consume more energy. Additionally, the faster the supercomputer, the more energy it consumes.
These observations are not rules, but are generally considered true in systems research. Even though the number of cores inside processors grows, more or less, exponentially, the single-core performance has relatively been steady the last decade. This means that the total cores is not far from describing the performance of the supercomputer (if we ignore accelerators). Similarly, although the lithography may have an influence on the efficiency of the CPU cores, it is usually the environment (i.e. heatsink, package size, TDP) that will dictate the final energy use of the supercomputer. Hence, there is the correlation between the total cores and energy.
More interestingly, if we take the colour mapping into account, we can see that the GPU-equipped supercomputers (yellow) consume less power than CPU-only supercomputers. This is expected, as the ranking of HPC facilities is based on the Flops/s (floating point operations per second) performance metric that is only helpful to certain types of problems [8]. This metric favours GPUs, because it assumes highly-parallel and highly-independent workloads, and with little control complexity, which is what CPUs are most appropriate for.
At the same time, it is also illustrated that both GPU-equipped and CPU-only supercomputers coexist in the rankings in a uniform way. This is an interesting observation and it holds because those CPUs are still optimised for HPC workloads, such as by having high core-counts per socket. This essentially means that the line between CPUs and GPUs in supercomputers is fading for CPU-only entries, as with rank 4 (uppermost purple point), where Fugaku’s A64FX processor has 52 cores (4 are helper cores) and a high-bandwidth memory (HBM2) [9], a feature usually found in accelerators.
Finally, if we consider the point-size dimension (core-counts per CPU), we can see that the GPU-equipped supercomputers generally feature smaller CPUs than CPU-only supercomputers. This makes perfect sense based on the latter insight that CPU-only supercomputers migrate some of the GPU functionality to the CPU sockets, such as by having a high core count (notably the higher ranking ones, which are mostly more recent). The big outlier in this figure is Sunway TaihuLight in China (largest point), the 13th supercomputer in this edition of the list, whose manycore CPUs combine 260 cores in a single socket [10].
In conclusion, this use case demonstrates that many-dimensional datasets can quickly be explored using the proposed framework, while helping to derive advanced and specialised insights from a single plot.
III-B World happiness report
Another interesting and timely dataset is the world happiness report for 2024, as served by Kaggle [11]. The idea behind this dataset is to use easily obtainable variables to calculate a ladder score; a research-informed model for happiness [12]. The original source is in the public domain (as with the GDP data from the World Bank). The study is conducted yearly by a partnership of reputable organisations including the Oxford Wellbeing Research Centre, with the aim to raise awareness for underperforming indicators per country. Hence, there are around 140 available multi-dimensional data-points (plus 3 with missing values), one for each country.
Figure 5: Multiple factors of world happiness. Point size is the happiness. The label filter facility (top) is used to decide which countries to annotate.
Figure 5 presents the generated plot, with additional visual cues for depth perception (a field of view of 70, slight fog, and depth-mapped point colouring). The axes dimensions are the variables health life expectancy, freedom to make life choices, the logarithm of GDP per capita (as given in the dataset). The 4th dimension is the point size, which represents the ladder happiness score. As can be seen by the 2 projections in the corresponding surfaces, there is a clear correlation between the $x$, $y$ and $z$ variables, before talking about the happiness index. The health life expectancy is linearly related to the Log GDP per capita, as seen from the linear trend in the projections of the right surface. In the other take-2 combinations of the 3 axes, we see a more clustered representation towards the high-end, where the points are also larger. This means that they need to co-exist to contribute to happiness, but are related less directly.
The 5th dimension is the country names, and is here partly represented by labels as a discrete dataset. In figure 5 (top) there is an overlaid screenshot from the control panel on the 6th dimension (perceived generosity) with which the label subset is based on. In this example, the top 9 countries based on the generosity variable are denoted with labels. As an observation, there is some correlation of happiness with generosity, but is much less influential than the other indicators. For instance, the points of Ethiopia and Kenya are in the top few with regard to generosity, but their medium size and central positioning demonstrates weaknesses in the other happiness indicators and the overall happiness score.
IV Empirical user study
The study is conducted online with the aim to identify and quantify the effectiveness of different visual cues in 3D Cartesian scatterplots. This is done by measuring the accuracy of the provided readings as perceived by participants and the time needed to answer them. The invited participants click on the study link in an email to answer 6 questions of the style “what is the $[X,Y,Z]$-coordinate of the marked data point?”. These 6 questions appear on 18 pages, 3 for every one of the axes, to simplify the presented instructions. The number 6 is selected based on the expected time to complete the quiz (see sections IV-A and VI-B). The plotted data follow a random uniform distribution per axis, resulting in discrete data. By doing so, a worst case is attempted with respect to any data localities in the dataset that may serve as additional visual cues. Figure 6 presents an example screenshot from the website during a test session.
Figure 6: Screenshot of a question from a quiz instance.
Figure 7: Image example per question type, using the same dataset and requesting the $y$-axis coordinate of the same highlighted point.
IV-A Question sampling
The study is designed to be informative yet short and inviting, hence there are 11 question types in total, taking up to around 5 minutes to complete. Each question type is a feature combination, and the rationale behind these combinations is to be incremental whenever is deemed appropriate. For instance, the three evaluated values of the field of view do not multiply the rest of our design space by three. The current findings could still be used more iteratively or exhaustively in future explorations (as in section VI-B). Table II summarises the question types in terms of visual cue combinations. Figure 7 shows an example of each question type using the same random dataset.
Table II: Visual cue combinations in the study
{NiceTabular}
¿c c ¿c c ¿l c ¿l[colortbl-like] Question&FoV ()FogColourProjectionsSurface ticksLight path
@@ IV-A Question sampling
1 The point colouring is here used to identify the point in question.
2 The point colouring is here used as a cue for depth perception.
@@
This number of possible question type combinations is (${11\choose 6}=462$, since each user gets a random subset of the 11 question types of size 6. Every question appears with a unique dataset to the others, and a random permutation of the datasets increases the uniqueness of each quiz instance with respect to the dataset order. After considering the $6!=720$ dataset permutations, we conclude that there are $720\times 462=332,640$ possible quiz instances, with varying amounts of similarity. The most important feature here is the uniqueness of the datasets between the questions of the same instance, to avoid unintentional knowledge transfer between different questions. Finally, considering multiple datasets per question is crucial for disassociating any biases related to the difficulty of the dataset (elaborated in section IV-D).
IV-B Software infrastructure

### v3

# Empirical Study on the Use of 3D Scatterplots as 2D Figures

@@ Abstract
3D scatterplots are a well-established plotting technique that can be used to represent data with three or more dimensions. On paper and computer monitors they are essentially two-dimensional projections of the three-dimensional Cartesian coordinate system. This transition from the 3D space to two dimensions is not done consistently among scientific software, as there is currently limited quantifiable evidence on the effectiveness of each approach. Notably, the frequent lack of visual cues such as with regard to depth perception is equivalent to a reduction of dimensionality by one. Hence, their use in manuscripts is less common or straightforward. In this empirical study, an online survey is conducted within an academic institution to identify and quantify the effectiveness of feature or feature combinations on 3D scatterplots in terms of reading time and accuracy.
keywords:
3D scatterplots, Cartesian, visual cues, depth perception, discrete datasets, survey
1 Introduction
@@
Effectively visualising and analysing multi-dimensional data is of prime importance for a variety of sciences, and is an invaluable tool in data mining, machine learning and statistics. With global trends and events such as large-language models [1] and pandemics [2], intuitively visualising large or complex data are already concerning a wider demographic. A visualisation method that is appropriate for multi-dimensional data is 3D scatterplots, as they can translate data for three or more dimensions into the 3D space, but extra care is needed to produce effective and informative plots.
@@
When compared to 2D scatterplots, having an extra dimension in space can be a blessing and a curse. It is a blessing because a single 3D scatterplot can give more advanced insights including how each of the dimensions correlate with each other or collectively. When combined with additional cues such as the point size, 3D scatterplots can also provide quicker insights for complex problems, especially when the data can be explained effectively with spatial information. It is also a curse, however, because in their static form as in scientific manuscripts, they need to be flattened to fit in two-dimensional means. These essentially reduce to 2D projections of (pseudo) 3D models. Such projections usually lack realism and visual cues to help with depth perception, for example, and this introduces ambiguity in the readings.
@@
This paper explores the hypothesis that the use of certain visual cues can significantly help with the readability and effectiveness of 3D scatterplots. The list of explored visual cues is the field of view, fog, depth-based colour mapping, projections, light paths (more broadly called drop lines) and surface ticks (elaborated in section 3). We evaluate the effectiveness of these visual cues using random data with a survey of 57 participants from an academic community. The survey measures the time and accuracy of the user measurements to identify impactful visual cue combinations and settings.
@@
As a motivation, figure 1 shows 3D scatterplots from popular open source plotting software (Gnuplot and GNU Octave) that demonstrate common weaknesses. The default configurations (without additional scripting or code) yield same-sized points, isometric axes, and absence of additional indicators to track the coordinates of the points. This experiment uses a random distribution dataset as a worse case (as in section 3.1), but the discrete nature of the data is easily relatable to existing problems. These plots can be considered counterexamples for effectively plotting 3D scatterplots, though more elaborate design decisions and software are capable of adding more realism [3].
@@
Figure 1: Default 3D plots from modern software on an arbitrary dataset
@@
The contribution of this paper is a systematic empirical user study that evaluates the a number of visual cues in a quantifiable way in terms of measurement error (accuracy) and time (also relating to intuitiveness). The visual cues are organised in question types (see section 3.1). The outcomes of this study can be used to increase the effectiveness of 3D scatterplots and plotting software by introducing any missing beneficial features, or enabling them by default. The identified gap in research is the absence of such a quantitative study, which can be the reason for software of highly-varying quality in 3D scatterplots, and their relatively limited use on paper. A series of hypotheses relating to the readability of 3D scatterplots are investigated:
  • H1.
    An exaggerated field of view (FoV) can be used to improve the readability of 3D scatterplots by improving depth perception.
  • H2.
    Adding fog to a 3D scatterplot can improve the readability of 3D scatterplots by improving depth perception.
  • H3.
    Applying a colour mapping corresponding to the distance from the object to the camera (observer) can improve the readability of 3D scatterplots by improving depth perception.
  • H4.
    Adding projections to the points that correspond to their size and shape can improve their readability by being able to distinguish them easier on the corresponding grid lines.
  • H5.
    A 2D version of the grid lines working as semi-transparent surface ticks can be used to improve point readability by improving spatial perception in at least the axis of the surface ticks.
  • H6.
    Adding light paths (similar to drop lines) to the points that correspond to their size and shape can improve readability by enhancing the association of projections to points.
  • H7.
    Certain combinations of the above can improve the readability of the points by complementing each other such as with respect to providing visual cues about different dimensions.
Following is a literature review on the area (section 2), the methodology of the survey (section 3), results (section 4), and finally a discussion (section 5) on the hypotheses and limitations, and a conclusion (section 6).
2 Related work
3 Methodology
The study is conducted online with the aim to identify and quantify the effectiveness of different visual cues in 3D Cartesian scatterplots. This is done by measuring the accuracy of the provided readings as perceived by participants and the time needed to answer them. The invited participants click on the study link in an email to answer 6 questions of the style “what is the $[X,Y,Z]$-coordinate of the marked data point?”. These 6 questions appear on 18 pages, 3 for every one of the axes, to simplify the presented instructions. The number 6 is selected based on the expected time to complete the quiz (see sections 3.1 and 5.2). The plotted data follow a random uniform distribution per axis, resulting in discrete data. By doing so, a worst case is attempted with respect to any data localities in the dataset that may serve as additional visual cues. Figure 2 presents an example screenshot from the website during a test session.
Figure 2: Screenshot of a survey question for the $x$ value of a point. Note that the questions come in triples, one for each $x$, $y$ and $z$ coordinate.
Figure 3: Image example per question type, using the same dataset and requesting the $y$-coordinate of the same highlighted point. Note that all $x$, $y$ and $z$ coordinates are asked by the survey.
3.1 Question sampling
The survey is designed to be informative yet short and inviting, hence there are 11 question types in total, taking up to around 5 minutes to complete. Each question type is a feature combination, and the design space exploration is not exhaustive. For instance, the three evaluated values of the field of view (FoV) in questions A, B and C do not multiply the rest of our design space by three, as a default is used as a baseline for other questions. A future study could utilise the presented results to build new explorations that focus on specific aspects of this study and new combinations of visual cues (see section 5.2). Table 1 summarises the question types in terms of visual cue combinations. Figure 3 shows an example of each question type using the same random dataset.
Table 1: Visual cue combinations in the survey
@@ 3.1 Question sampling
>c c >c c >l c >l[colortbl-like] Question&FoV ()FogColourProjectionsSurface ticksLight path
@@
1 The point colouring here is used to identify the point in question.
2 The point colouring here is used as a visual cue for depth perception.
The main rationale behind this selection of questions is to be able to address the hypotheses as introduced in section 1. Hypothesis H1 on FoV is investigated by the combination of question types (A, B, C), hypothesis H2 on fog by the combination (B, D), hypothesis H3 on depth-based colour-mapping by questions (B, E) and hypothesis H5 on surface ticks by questions (H, I). Hypothesis H4 on projections is mainly explored by (B, F), and H6 on light paths is explored incrementally with (B, H, J, K) corresponding to 0, 1, 2 and 3 light paths. Hypothesis H7 on combinations is partly addressed by questions with overlapping features such as G that enables colour-mapping and projections, or by the incremental aspect of some of the question combinations like (B, H, J, K) for light paths, as well as their assumed coexistence with their corresponding projections. See section 5.1 for the final discussion on the hypotheses.
@@
This number of possible question type combinations is (${11\choose 6}=462$, since each user gets a random subset of the 11 question types of size 6. Every question appears with a unique dataset to the others, and a random permutation of the datasets increases the uniqueness of each quiz instance with respect to the dataset order. After considering the $6!=720$ dataset permutations, we conclude that there are $720\times 462=332,640$ possible quiz instances, with varying amounts of similarity. The most important feature here is the uniqueness of the datasets between the questions of the same instance, to avoid unintentional knowledge transfer between different questions. Finally, considering multiple datasets per question is crucial for disassociating any biases related to the difficulty of the dataset (elaborated in section 4).
3.2 Software infrastructure